OpenStudy (anonymous):

Divide and Simplify: y^2-25 over y/ y-5 over y-2.

OpenStudy (anonymous):

Rearranging (y^2-25)(y-2) / y(y-5) = (y+5)(y-2) /y = y +3 - (10/y)